# Exploring and Evaluating Real-world CXL: Use Cases and System Adoption

### v1

@@ Abstract.
Compute eXpress Link (CXL) is emerging as a promising memory interface technology. Because of the common unavailiability of CXL devices, the performance of the CXL memory is largely unknown. What are the use cases for the CXL memory? What are the impacts of the CXL memory on application performance? How to use the CXL memory in combination with existing memory components? In this work, we study the performance of three genuine CXL memory-expansion cards from different vendors. We characterize the basic performance of the CXL memory, study how HPC applications and large language models can benefit from the CXL memory, and study the interplay between memory tiering and page interleaving. We also propose a novel data object-level interleaving policy to match the interleaving policy with memory access patterns. We reveal the challenges and opportunities of using the CXL memory.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Compute eXpress Link (CXL) is emerging as a promising memory interface technology. Based on the standard PCIe serial interface, CXL can attach memory to the CPU and appears to a system as a CPU-less NUMA node. This introduces flexibility in memory subsystem design and fine-grained control over memory bandwidth and capacity.
Compared with the traditional memory interface technology (such as DDR), CXL provides multiple benefits. First, CXL provides a more scalable method than DDR to increase memory capacity and bandwidth, because CXL does not increase the number of DDR channels or cause super-linear increase of energy consumption and signaling challenges as DDR (Sun et al., 2023). Second, the CXL memory provides a solution to avoid memory stranding (Li et al., 2023), because using the CXL memory as a pool, the stranded memory can be disaggregated from servers and provide memory capacity as demanded. Third, with CXL, the memory technologies are decoupled from the memory interface technologies used by the CPU (Wang et al., 2020; Xiang et al., 2022; Yang et al., 2020), which provides unprecedented flexibility to reuse old memory and build the memory pool. Fourth, CXL allows the CPU to communicate with the attached memory in a cache-coherent fashion using load and store instructions. This is aligned with the traditional memory semantics, and hence provides conveniences of using existing programming models and system designs (e.g., NUMA-based tiered memory). Fifth, data fetched from and written to the CXL memory can be cached in the cache hierarchy in the CPU, which allow the system to explore locality for higher performance. This contrasts with RDMA where data cannot be cached in the CPU.
@@ 1. Introduction
However, CXL memory introduces longer memory access latency. This longer latency comes from PCIe, CXL memory controller, and CXL home agent (HA) on the CPU. Figure 1 compares local NUMA, traditional remote NUMA, and CXL-based memory expansion in terms of memory latency. It has been reported that the CXL adds about 70-150 $ns$ of extra latency over normal DRAM access (Maruf et al., 2023).
Figure 1. Breakdown of CXL memory access latency.
@@
Given the CXL performance, we face a series of questions: what are the use cases for the CXL memory? What are the impacts of real CXL memory on application performance? At the application level, how to use the CXL memory in combination with fast memory components (e.g., using uniform page-level interleaving vs. data object-level interleaving vs. memory binding)? This paper aims to discuss those questions, characterize the performance of the CXL memory, and explore various paths to use the CXL memory. We study three genuine CXL memory expansion cards from different vendors, and have the following insights.
@@
The CXL memory is a unique “NUMA node”. The CXL memory appears as a CPU-less NUMA node. In the three systems we evaluate, the CXL memory appears as a two-hop-away NUMA node in terms of access latency. Depending on the memory vendors, the peak bandwidth of the CXL memory varies a lot, ranging from 9.8% to 80.3% of the peak bandwidth of local DRAM. In addition, as we scale the number of threads to access the CXL memory, its memory bandwidth is easily saturated because of PCIe (the saturation point is reached when the number of threads is only 4), while local DRAM (or LDRAM for short) and remote DRAM-based NUMA node (or RDRAM for short) have much better scalability. This scaling difference between the CXL memory and DRAM highlights the importance of appropriately distributing memory accesses between them for high performance. Also, when the system is under heavy load, we notice that LDRAM and RDRAM can show the similar latency as the CXL memory because of the contention on the memory controller (MC) or data path, which shows the potential of using CXL as LDRAM or RDRAM for latency-sensitive applications when the system is under heavy load.
HPC applications can benefit a lot from the CXL memory for high performance while saving fast memory. Our work goes beyond the existing work (Sun et al., 2023) that focuses on applications exhibiting $ms$-scale latency (e.g., social network microservices) to study the potential of the CXL memory. We study a spectrum of HPC workloads, covering the most common and representative “HPC dwarfs” (Asanovic et al., 2006). We reveal that some HPC applications (such as CG and BT (Löff et al., 2021)) can tolerate low bandwidth and high latency of the CXL memory under certain scales (the performance loss is less than 3.2%, compared with LDRAM), because of their compute-intensive nature.
In addition, the page interleaving policy, embraced by the industry (e.g., Micron and AMD (Petrucci et al., 2024), Astera Labs (Labs, [n. d.]), and Samsung (Prout, [n. d.])) as an application-transparent technique to integrate CXL with the existing memory components, provides opportunities to save LDRAM for HPC applications. When interleaving CXL and RDRAM, we see minor performance difference from interleaving CXL and LDRAM for some applications. This is because the CXL memory dominates the memory performance, and the performance of other memory components has minor impact on the overall performance.
To maximize the interleaving performance, we propose a data object-level interleaving policy. This policy decides whether memory pages allocated to a data object should be interleaved between CXL and DRAM or allocated to LDRAM first (“LDRAM preferred”). This policy maximizes memory bandwidth (or minimize latency) for data objects whose accesses favor high bandwidth (or low latency). This policy reduces the usage of fast memory (LDRAM) by $48\%$ on average and outperforms the uniform interleaving policy (the Linux policy) by $65\%$ on average.
@@
Using the CXL memory for large language models (LLM) faces challenges. We study the cases that use the CPU memory as an extension to the GPU memory to enable LLM training (using ZeRO-Offload (Ren et al., 2021c)) and inference (using FlexGen (Sheng et al., 2023)) such that we can use less GPUs for LLM without the constraint of GPU memory capacity. This method has to frequently copy tensors between the GPU and the CPU, and offloads certain computations from the CPU to the GPU to maximize GPU memory saving or reduce I/O offload. This method (named tensor offloading in the rest of the paper) has been commonly studied and deployed in industry.
@@
We find that using tensor offloading based upon the CXL memory brings limited performance improvement. This is because the tensor copy between the CXL memory and GPU memory goes through a longer data path than memory accesses directly from CPU. Using CXL 1.1 on our platform, this data path is bottlenecked by the PCIe interconnect between CPU and GPU. As a result, adding CXL cannot show the bandwidth benefits for tensor offloading. In contrast, the computation offloaded to the CPU (i.e., the optimizer for LLM training and the attention computation for LLM inference), which tends to be bandwidth sensitive, can benefit from the extra bandwidth of CXL. In addition, the CXL memory increases memory capacity and allows us to use larger batch sizes for LLM inference, leading to throughput gains.
@@
Memory tiering solutions need to be improved. Treating the CXL memory as a memory tier, existing work (Corbet, [n. d.]; Verma., 2022; Ren et al., 2024; Maruf et al., 2023; Raybuck et al., 2021; Lee et al., 2023) migrates pages between the CXL memory and fast memories based on page access frequency or recency (i.e., hotness). Those solutions are seldom studied with the real CXL memory, and how they interplay with the existing system (e.g., page interleaving) is largely unknown.
@@
We find that the dynamic page migration in the memory tiering solutions are not integrated well with the static page interleaving, because of invalidness of NUMA hint faults. Depending on the temporal and spatial distribution of hot pages, the dynamic page migration can degrade application performance compared to no migration. We also observe that the old-fashioned NUMA first touch and Tiering-0.8 (Verma., 2022) (the most recent Linux Patch for AutoNUMA to support memory tiering) is very effective, outperforming a set of page migration and interleaving solutions.
2. Background
2.1. Compute Express Link
The CXL specification defines three protocols: CXL.io for initial discovery and configuration of CXL devices, CXL.cache for CXL-devices to access the host memory using PCIe protocol features, and CXL.mem for the host access to CXL device memory. The Home Agent (HA) on the CPU and CXL controller on the device manage the interaction between the host and CXL device. The HA administers the CXL.mem protocol and seamlessly presents the CXL memory to the host as memory on a remote NUMA node.
@@ 2.1. Compute Express Link
There are three types of CXL devices. The type-3 device is related to our evaluation. Such a device supports CXL.io and CXL.mem, and is used for memory bandwidth or capacity expansion, and storage class memory in memory tiering.
CXL.mem provides three coherence models for CXL exposed host-managed device memory (HDM): HDM-H (host-only coherent and used for type-3), HDM-D (device coherent relying on CXL.cache to manage coherence with the host and used for type-2), and HDM-DB (device coherent using back-invalidation and used for type-2 or type-3).
@@
The CXL specification has been going through three major versions: 1.1, 2.0, and 3.0. 1.1 focuses on directly-attached CXL devices, 2.0 incorporates switch-based pooling, and 3.0 supports switch-less pooling and higher bandwidth.
@@
Most of the real CXL devices nowadays are HDM-H using CXL 1.1. The three devices for our evaluation are among them.
Table 1. Three systems with CXL devices.
@@
CXL B Single channel DDR5, memory
@@ 2.2. CXL Systems for Evaluation
CXL requires compatible hardware in both the CPUs and peripheral devices. Notably, the 4th-generation Intel Xeon Scalable Processors (such as Sapphire Rapids) and the 4th-generation AMD EPYC Processors (such as Genoa) are among the first mainstream server CPUs to support the CXL 1.1. Several CXL memory devices have been developed as commercial products by leading hardware manufacturers such as SK hynix, Astera Labs, Montage, Samsung, and Micron. We use three CXL devices from three vendors. Table 1 summarizes the three systems equipped with the devices.
@@
The system A listed in Table 1 has two sockets (0 and 1), and a CXL device is attached to the socket 1 by a CXL link over PCIe 5.0. This system has two AMD EPYC 9354 CPUs. Accessing the CXL memory from CPU 0 has to go through the HyperTransport interconnect, which leads to longer latency than accessing from CPU 1. Different from the traditional NUMA node, the CXL memory device lacks CPU cores and caches, and does not have an expensive interconnect between the CXL IP and the device’s memory controller. Hence, the CXL memory is represented as a CPU-less NUMA node in the system A. From the view of a CPU, there are three NUMA nodes: local DDR (or LDRAM), remote DDR on the other socket (RDRAM), and CXL memory. The system B has the same organization as the system A. The system C has a slightly different organization: the CXL device is attached to the socket 0 (instead of socket 1).
3. Basic Performance Characteristics
@@
Figure 2. Load latency with random and sequential memory accesses to a cache block.
@@
Evaluation methodology. We evaluate memory latency and bandwidth using Intel Memory Latency Checker (MLC) (Intel Corporation, 2019). MLC disables hardware prefetcher for Intel processors (the systems B and C), but cannot do so for AMD processors (the system A). For latency tests, MLC uses typical pointer chasing. For each latency test, we repeat the test by 5,000 times, and report the average value after excluding outliers (caused by operating system services and random TLB misses). For bandwidth tests, we use MLC to perform sequential and random memory accesses. The sequential accesses in combination with thread-level parallelism introduces parallel memory accesses, revealing peak memory bandwidth. For each bandwidth test, we repeat the test by 2,000 times, and report the average value.
@@
Figure 3. Bandwidth scaling for data loading.
Latency results. See Figure 2. We have two observations. Basic observation 1. Compared with LDRAM, CXL is much slower than one-hop-away NUMA node (RDRAM). For example, for sequential accesses, compared to LDRAM, CXL increases the latency by 2.8$\times$, 2.4$\times$, and 2.2$\times$ on the three systems respectively, while one-hop-away NUMA increases the latency by 1.8$\times$, 1.7$\times$, and 1.5$\times$. In fact, assuming that adding a hop of NUMA distance introduces a constant latency in a system, the CXL memory is comparable to a two-hop-away NUMA node, in terms of access latency.
@@
Basic observation 2. Figure 2 shows that the CXL memory from different vendors show quite different latency. For example, for sequential accesses, the CXL memory in the system A adds latency by 153 $ns$, while the CXL memory in the system B adds latency by 211 $ns$, compared to LDRAM. Since the two systems use the same PCIe and DRAM technologies, such a latency difference mainly comes from the difference in the CXL controller and HA on the CPU.
Figure 4. Memory latency and bandwidth under varying load.
Bandwidth results. Figure 3 shows the bandwidth scaling as we change the number of threads. We have three observations.
@@
Basic observation 3. The CXL memory bandwidth is saturated as the number of threads is over 8, while the saturation points for LDRAM and RDRAM are much higher than that in CXL (e.g., 28 and 20 on the system B), indicating that the CXL memory bandwidth is limited by PCIe.
Basic observation 4. The peak CXL memory bandwidth is much lower than that of LDRAM (the CXL memory bandwidth is 9.8%, 28.1%, and 80.3% of LDRAM bandwidth on the systems A, B, and C respectively). The peak CXL memory bandwidth is also lower than that of RDRAM (the CXL memory bandwidth is 17.1% and 46.4% of the RDARM bandwidth on the systems A and B respectively), but can be pretty close to the RDRAM bandwidth (see the system C).
@@
Basic observation 5. Difference in bandwidth scaling between LDRAM, RDRAM, and CXL highlights the importance of distributing memory accesses between them. For example, in the system B, to maximize the bandwidth usage, we would assign 6, 23, and 23 threads to access CXL, LDRAM, and RDRAM respectively. This leads to a peak bandwidth of 420 GB/s, larger than any other thread assignment.
@@
Basic observation 6. The latency of accessing LDRAM and RDRAM can be similar to the CXL memory when reaching their peak memory bandwidth. For example, in Figure 4.(c), once the bandwidth consumption of LDRAM and RDRAM approaches the peak (110GB/s for LDRAM and 84GB/s for RDRAM), their latencies are as high as 543 $ns$ and 600 $ns$ respectively, which are pretty close to the CXL memory latency (400 $ns$ 550 $ns$) when approaching the peak bandwidth of the CXL memory under heavy load. This demonstrates the potential of using CXL as LDRAM and RDRAM when the system is under heavy load.
@@
Takeaway: CXL memory performs as a NUMA node with latency similar to RDRAM but much lower (or comparable) bandwidth. But different from the traditional NUMA node, the CXL memory is unique in terms of performance scalability and performance under heavy load.
4. HPC Workloads
Figure 5. The overall performance for all benchmarks using various memory components. “XS” stands for “XSBench”.
We analyze HPC workload performance to demonstrate the use case of CXL memory, focusing on seven workloads from NPB benchmark suite (Löff et al., 2021) and XSBench benchmark (Tramm et al., 2014). These workloads, which require processing a large working memory set using multiple threads, are summarized in Table 2. They collectively cover the most common and representative HPC applications (Asanovic et al., 2006). In our study, we use various memory allocation policies, including preferred and interleaving. The preferred policy for a specific memory node indicates that the memory is allocated in that memory node first; when that memory node runs out of space, the page allocation goes to another memory node closet to the CPU according to the NUMA distance. The interleaving policy indicates that pages are allocated among memory nodes in a round robin fashion. We present evaluation results on the system A, because its performance is between the performance of the systems B and C, hence can be representative.
Table 2. HPC workloads for evaluation. “BW” stands for “bandwidth”.
BT Dense linear algebra Unit-strided memory accesses from dense matrices Class E 166 GB u, rsh, forcing
LU Sparse linear algebra Indexed loads and stores from compressed matrices Class E 134 GB u, rsd
CG Sparse linear algebra Irregular memory accesses based on indirect indexing Class E 134 GB a
MG Structured grids Dynamic updates based on subdivided regular grids Class E 210 GB v, r
SP Structured grids Intense floating-point computations for linear equations Class E 174 GB u, rsh, forcing
FT Spectral method Bandwidth-consuming matrix transpose Class D 80 GB u0, u1
4.1. Evaluation Results
HPC observation 1. Using CXL preferred, some HPC applications show minor performance loss, compared to using LDRAM and RDRAM.
Figure 5 shows the performance. We observe that BT and LU, when the number of threads is small (4), the performance difference between LDRAM and CXL preferred is very small (less than 3.2%). These benchmarks are highly compute intensive and not sensitive to long memory latency and bandwidth with a small number of threads. These results show great potential of using CXL memory to save LDRAM. For other benchmarks, the performance difference can be up to 1.9×1.9$\times$1.9 × when the number of threads is 4.
We also notice that when the number of threads is large (32), the performance difference between LDRAM and CXL preferred is at least 1.9×1.9$\times$1.9 ×, and the performance difference between RDRAM and CXL preferred is at least $\times1$. When the number of threads is large, the memory-level parallelism imposes high pressure on memory components, exacerbating inferior performance of the CXL memory.
HPC observation 2. Bandwidth-sensitive and latency-sensi
Figure 6 shows the results for MG (bandwidth-sensitive) and CG (latency-sensitive). For MG, we see that when the number of threads increases from 4 to 32, the performance of “interleave all” (i.e., interleaving between LDRAM, RDRAM and CXL, which achieves the highest bandwidth) is consistently better than that of CXL preferred by 10%-85%. Meanwhile for CG, the performance of “interleave all” performs worse than that of CXL preferred by up to 1.6$\times$, because using CXL preferred, consecutive memory accesses tend to fall into the same memory node, which is favored by CG for high performance (see more discussions for HPC observation 4).
Figure 6. The scalability study for CG and MG. The performance is normalized by that of using LDRAM only.
HPC observation 3. When the interleaving policy involves the CXL memory, we can save local fast memory by using remote fast memory.
Figure 7 shows the results with various interleaving policies, and the benchmarks run on CPU 0. We can see that the performance difference between the interleaving RDRAM+CXL and interleaving LDRAM+CXL is less than $9.2\%$ for all benchmark. In other words, interleaving CXL and RDRAM, we do not need local fast memory to achieve a similar performance.
The above results are because of large performance gap between the CXL memory and DDR (e.g., 2.1×2.1$\times$2.1 × and 1.2×1.2$\times$1.2 × longer in memory access latency, compared to LDRAM and RDRAM respectively.): because of data dependency and limited hardware resources (e.g., MSHR and the queues in MC), the performance is highly impacted by the slow CXL memory and irrelevant whether the LDRAM or RDRAM is utilized.
Figure 7. The performance of various interleaving policies for HPC applications.
HPC observation 4. The CXL memory can show unexpected high performance for the latency-sensitive applications with random memory accesses.
The CG results in Figures 6 and  7 support the above observation. CG is a latency-sensitive application because of its indirect indexing-based memory accesses. In Figure 6, CXL preferred performs better than RDRAM preferred by $10.9\%$-$57.2\%$ when the number of threads is 4-20. This seems to be counter-intuitive, because the CXL memory has higher latency than RDRAM. We think that this superior performance may come from the optimization in the CXL device or customized caching policy in the processor for expensive, CPU-less memory accesses (Sun et al., 2023), and this optimization is especially effective for CG-style workloads. When the number of threads is larger than 20, the CXL memory accesses become more intensive, and inferior performance of the CXL memory becomes more obvious.
4.2. Object-Level Interleaving
Instead of generally interleaving pages at the application level (named application-level interleaving), we propose a method to interleave pages at the data object level. This enables fine-grained control over how pages are interleaved. Since different data objects have different access patterns, using the fine-grained control allows the bandwidth-sensitive object to be accessed with high bandwidth, while the latency-sensitive object is allocated locally for high performance.
Interleaving method. We use two criteria to select data objects and allocate them using the interleaving policy.
With high thread-level parallelism, memory accesses to the above data objects are more sensitive to memory bandwidth. Besides the above data objects, the memory allocation for other objects uses the “preferred” policy. The last column in Table 2 summarizes those bandwidth-hungry data objects selected for interleaving.
We evaluate the effectiveness of our object-level interleaving using the following approach. In each test, we run the workload on CPU 0 using both LDRAM (memory node 0) and CXL memory. The LDRAM is either 64GB or 128GB, and the memory consumption of all applications exceeding 64GB, which allows us to assess the cases with both sufficient and insufficient LDRAM. The CXL memory is consistently 128GB. We make the following object-level interleaving observations (abbreviated as OLI observations).
Figure 8. Performance speedup of various memory allocation strategies compared to “LDRAM preferred”, using sufficient LDRAM (128GB).
OLI observation 1. When the local DRAM is sufficient (with 128 GB), the object-level interleaving performance is similar to that with LDRAM preferred, and consistently outperforms the application-level interleaving in all scenarios.
Figure 8 shows that OLI performance is 65% better than the application-level interleaving on average, demonstrating the effectiveness of OLI to meet the diverse requirements of latency-sensitive and bandwidth-sensitive objects. Additionally, we observe that by using OLI, we can effectively reduce the fast memory size by an average of 48% (and up to 80% for FT), while still achieving similar performance as LDRAM preferred. The performance difference between OLI and LDRAM preferred is less than 1% on average for all HPC workloads, except XSBench, when LDRAM is sufficient.
Figure 9. Performance speedup of various memory allocation strategies, compared to “LDRAM preferred”, using insufficient LDRAM (64GB).
Figure 9 shows that on average, OLI performs 1.42$\times$ better than LDRAM preferred (and up to 2.35$\times$ for BT). It also outperforms the application-level interleaving by 1.32$\times$ on average (and up to 1.84$\times$ for LU). This is because of the following reasons. (1) The performance of LDRAM preferred is highly related to when the data objects are allocated. If LDRAM is full, latency-sensitive objects might end up in the slower CXL memory. (2) The bandwidth-sensitive data objects do not have opportunities to be allocated on both LDRAM and CXL for high bandwidth.
In XSBench, LDRAM preferred performs better than both application-level and object-level interleaving. This is because most of the memory accesses in XSBench are concentrated in a small, latency-sensitive memory set.
Takeaway: the CXL memory offers additional bandwidth, but HPC workloads may not benefit from it directly. The application-level interleaving can undermine performance, while the data object-level interleaving delivers performance comparable to or better than the default LDRAM-centric page allocation while saving fast memory size.
5. CXL for Large Language Models

### v2

@@ Abstract
Compute eXpress Link (CXL) is emerging as a promising memory interface technology. Because of the common unavailability of CXL devices, the performance of the CXL memory is largely unknown. What are the use cases for the CXL memory? What are the impacts of the CXL memory on application performance? How to use the CXL memory in combination with existing memory components? In this work, we study the performance of three genuine CXL memory-expansion cards from different vendors. We characterize the basic performance of the CXL memory, study how HPC applications and large language models can benefit from the CXL memory, and study the interplay between memory tiering and page interleaving. We also propose a novel data object-level interleaving policy to match the interleaving policy with memory access patterns. We reveal the challenges and opportunities of using the CXL memory.
footnotetext: *Co-first authors.
I Introduction
@@
Compute eXpress Link (CXL) is a promising memory interface technology. Based on the standard PCIe serial interface, CXL attaches memory to the CPU and appears as a CPU-less NUMA node. The CXL memory can be accessed in a cache-coherent fashion using load/store instructions. However, CXL memory introduces longer memory access latency. This longer latency comes from PCIe, CXL memory controller, and CXL home agent (HA) on the CPU. Figure 1 compares local NUMA, traditional remote NUMA, and CXL-based memory expansion in terms of memory latency.
@@
Given the CXL performance, we face a series of questions: what are the use cases for the CXL memory? What are the impacts of real CXL memory on application performance? At the application level, how to use the CXL memory in combination with fast memory components (e.g., using uniform page-level interleaving vs. data object-level interleaving vs. memory binding)? This paper aims to discuss those questions, and explore various paths to use the CXL memory. We study three genuine CXL memory expansion cards instead of using memory simulation or emulation.
@@
The CXL memory is a unique “NUMA node”. The CXL memory appears as a CPU-less NUMA node. In the three systems we evaluate, the CXL memory appears as a two-hop-away NUMA node in terms of access latency. Depending on the memory vendors, the peak bandwidth of the CXL memory varies a lot, ranging from 9.8% to 80.3% of the peak bandwidth of local DRAM. Moreover, as we increase the number of threads accessing the CXL memory, its bandwidth is quickly saturated due to the limited data transfer rate in CXL attached memory instead of PCIe bus — with saturation occurring when the number of threads reaches just four. In contrast, local DRAM (LDRAM) and remote DRAM-based NUMA nodes (RDRAM) exhibit much better scalability. This scaling difference between the CXL memory and DRAM highlights the importance of appropriately distributing memory accesses between them for high performance. Also, when the system is under heavy load, we observe that the latencies of LDRAM and RDRAM are similar to that of the CXL memory because of the contention on the memory controller (MC) or data path, which shows the potential of using CXL as LDRAM or RDRAM for latency-sensitive applications.
Figure 1: Breakdown of CXL memory access latency.
@@
Using the CXL memory for large language models (LLM) faces challenges. LLM can be memory-consuming and have execution phases sensitive to memory bandwidth [27, 54, 46], hence having potentials to benefit from CXL. We study the cases that use the CPU memory as an extension to the GPU memory to enable LLM training (using ZeRO-Offload [54]) and inference (using FlexGen [63]) such that we can use less GPUs for LLM without the constraint of GPU memory capacity. This method has to frequently copy tensors between the GPU and CPU, and offloads certain computations from the CPU to the GPU to maximize GPU memory saving or reduce I/O offload. This method (named tensor offloading in the rest of the paper) has been commonly studied and deployed in industry.
@@
We find that using tensor offloading based upon the CXL memory brings limited performance improvement. This is because the tensor copy between the CXL memory and GPU memory goes through a longer data path than memory accesses directly from CPU. Using CXL 1.1 on our platform, this data path is bottlenecked by the PCIe interconnect between CPU and GPU. As a result, adding CXL cannot show the bandwidth benefits for tensor offloading. In contrast, the computation offloaded to the CPU (i.e., the optimizer for LLM training and the attention computation for LLM inference), which tends to be bandwidth sensitive, can benefit from the extra bandwidth of CXL. In addition, the CXL memory increases memory capacity and allows us to use larger batch sizes for LLM inference, leading to throughput gains. Our study is unprecedented, because of its practicality of using real CXL for GPU-based LLM, different from CPU-based study [68].
CXL can be used to save fast memory without causing performance loss, and using application semantics to guide page interleaving for CXL can maximize CXL benefits. Our work goes beyond the existing work [67, 68] that focuses on applications exhibiting $ms$-scale latency (e.g., social network microservices) or commercial workloads to study the potential of the CXL memory. We study a spectrum of HPC workloads, covering the most common and representative “HPC dwarfs” [6]. We reveal that some HPC applications (such as CG and BT [15]) can tolerate the low bandwidth and high latency of the CXL memory under certain scales (the performance loss is less than 3.2%, compared with LDRAM), because of their compute-intensive nature.
In addition, the page interleaving policy, embraced by the industry (e.g., Micron and AMD [47], Astera Labs [30], and Samsung [48]) as an application-transparent technique to integrate CXL with the existing memory components, provides opportunities to save LDRAM for HPC applications. When interleaving CXL and RDRAM, we see minor performance difference from interleaving CXL and LDRAM for some applications. This is because the CXL memory dominates the memory performance, and the performance of other memory components has minor impact on the overall performance.
To maximize the interleaving performance, we introduce a novel data object-level interleaving policy. Different from Linux uniform page-level interleaving [67], this policy decides whether memory pages allocated to a data object should be interleaved between CXL and DRAM or allocated to LDRAM first (“LDRAM preferred”). This policy maximizes memory bandwidth (or minimizes latency) for data objects whose accesses favor high bandwidth (or low latency). This policy reduces LDRAM usage by $32\%$ and outperforms the uniform interleaving policy (Linux default) by $65\%$ on average.
@@
Memory tiering solutions need to be improved. Treating the CXL memory as a memory tier, existing work [13, 73, 55, 44, 50, 32] migrates pages between the CXL memory and fast memories based on page access frequency or recency (i.e., hotness). Those solutions are seldom studied with the real CXL memory, and how they interplay with the existing system (e.g., page interleaving) is largely unknown.
@@
We find that the dynamic page migration in memory tiering are not integrated well with the static page interleaving, because of invalidness of NUMA hint faults. Depending on temporal and spatial distribution of hot pages, the dynamic page migration can degrade performance compared to no migration. We also observe that the old-fashioned NUMA first touch and Tiering-0.8 [73] (the most recent Linux Patch for AutoNUMA to support memory tiering) is very effective, outperforming a set of page migration and interleaving solutions.
II Background
II-A Compute Express Link
@@
The CXL specification defines three protocols: CXL.io, CXL.cache, and CXL.mem. There are three types of CXL devices. The type-3 device is related to our evaluation. Such a device supports CXL.io and CXL.mem, and is used for memory bandwidth or capacity expansion in memory tiering.
@@
The CXL specification has been going through three major versions: 1.1, 2.0, and 3.0. CXL 1.1 focuses on directly-attached CXL devices, CXL 2.0 incorporates switch-based pooling, and 3.0 supports switch-less pooling and higher bandwidth.
@@
Most of the real CXL devices nowadays are host-managed device memory with host-only coherent (HDM-H) using CXL 1.1. The three devices for our evaluation are among them.
TABLE I: Three systems with CXL devices.
@@
max bandwidth 307.2 GB/s per socket
CXL B Single channel DDR5-8000, memory
@@ II-B CXL Systems for Evaluation
CXL requires compatible hardware in both the CPUs and peripheral devices. The 4th-generation Intel Xeon Scalable Processors (such as Sapphire Rapids) and the 4th-generation AMD EPYC Processors (such as Genoa) are among the first mainstream server CPUs to support the CXL 1.1. Several CXL memory devices have been developed as commercial products by leading hardware manufacturers such as Micron. We use three CXL devices from three vendors. See Table I.
@@
The system A in Table I has two sockets (0 and 1), with a CXL device attached to Socket 1 by CXL link over PCIe 5.0. Accessing CXL memory from Socket 0 goes through HyperTransport interconnect, leading to longer latency than accessing from Socket 1. From the view of a CPU, there are three NUMA nodes: local DDR (LDRAM), remote DDR on the other socket (RDRAM), and CXL memory. The system B has the same organization as A. The system C has a different organization: the CXL device is attached to Socket 0 (not 1).
III Basic Performance Characteristics
@@
Figure 2: Load latency with random and sequential accesses to a cache block.
@@
Evaluation methodology. We evaluate memory latency and bandwidth using Intel Memory Latency Checker (MLC) [22]. MLC disables hardware prefetcher for Intel processors (the systems B and C), but cannot do so for AMD processors (the system A). For latency tests, MLC uses typical pointer chasing. For each latency test, we repeat the test 5,000 times, and report the average value after excluding outliers (caused by operating system services and random TLB misses). For bandwidth tests, we use MLC to perform sequential and random memory accesses. The sequential accesses in combination with thread-level parallelism introduces parallel memory accesses, revealing peak memory bandwidth. For each bandwidth test, we repeat the test 2,000 times, and report the average value.
@@
Figure 3: Bandwidth scaling for data loading. Note that the scales of the figures for the three systems are different.
@@
Latency results. See Figure 2. Compared with LDRAM, CXL is much slower than one-hop-away NUMA node (RDRAM). In fact, assuming that adding a hop of NUMA distance introduces a constant latency in a system, the CXL memory is comparable to a two-hop-away NUMA node, in terms of access latency. Figure 2 shows that the CXL memory from different vendors show quite different latency. For example, for sequential accesses, the CXL memory in the system A adds latency by 153 $ns$, while the CXL memory in the system B adds latency by 211 $ns$, compared to LDRAM. Since the two systems use the same PCIe and DRAM technologies, such a latency difference mainly comes from the difference in the CXL controller and HA on the CPU.
Figure 4: Memory latency and bandwidth under varying load.
@@
Bandwidth results. Figure 3 shows that bandwidth scaling of LDRAM, RDRAM, and CXL are different as we change the number of threads. The CXL memory bandwidth is saturated as the number of threads is over 8, while the saturation points for LDRAM and RDRAM are much higher than that in CXL (e.g., 28 and 20 on the system B), because the bandwidth of CXL memory is constrained by the bandwidth of a single DDR channel. The peak CXL memory bandwidth is lower than that of RDRAM (the CXL memory bandwidth is 17.1% and 46.4% of the RDRAM bandwidth on the systems A and B respectively), but can be close to the RDRAM bandwidth (see the system C).
@@
The difference in bandwidth scaling between LDRAM, RDRAM, and CXL highlights the importance of distributing memory accesses between them. For example, in the system B, to maximize the bandwidth usage, we would assign 6, 23, and 23 threads to access CXL, LDRAM, and RDRAM respectively, because increasing thread counts beyond these points does not improve the bandwidth, shown in Figure 3(d). Using the above thread counts can lead to a peak bandwidth of 420 GB/s, larger than any other thread assignment.
@@
Basic observation. The latency of accessing LDRAM and RDRAM can be similar to the CXL memory when reaching their peak memory bandwidth. For example, in Figure 4(c), once the bandwidth consumption of LDRAM and RDRAM approaches the peak (110GB/s for LDRAM and 84GB/s for RDRAM), their latencies are as high as 543 $ns$ and 600 $ns$ respectively, which are pretty close to the CXL memory latency (400 $ns$ 550 $ns$) when approaching the peak CXL bandwidth under heavy load. This demonstrates the potential of using CXL as LDRAM and RDRAM under heavy load.
@@
Takeaway: CXL memory performs as a NUMA node with latency similar to RDRAM but lower (or comparable) bandwidth. Different from the traditional NUMA node, the CXL memory is unique in terms of performance scalability and performance under heavy load.
IV CXL for Large Language Models